OpenStudy (annie96):

help. Write the expression in simplified radical form. Show all work. 2sqrt(24)+4sqrt(6) Write the expression in simplified radical form. Show all work. (3/sqrt(5)) Solve 2 ^3sqrt(2x+1)=6 Given the complex number -4 - 2i, a.) Graph the complex number in the complex plane. b.) Calculate the modulus. Put answer in simplified radical form. Show all work.

OpenStudy (annie96):

graph to number 4

OpenStudy (anonymous):

2 \sqrt{24}+4 \sqrt{6}=4 \sqrt{6}+4 \sqrt{6}=8 \sqrt{6}

OpenStudy (annie96):